OpenStudy (anonymous):

M(2,4) is the midpoint of RS. If S has coordinates (9,11) , find the coordinates of R.

OpenStudy (angela210793):

XM=(XR+XS)/2 and YM=(YR+YS)/2 so 2=(XR+9)/2 and 4=(YR+11)/2 XR=4-9 and YR=8-11 XR=-5 and YR=-3 so R(-5:-3)